one positive number is 5 larger than another number. the sum of their squares is 53. find the numbers.
Answer by Jk22(389)

Let the number be a,b, the larger a=5+b

sum of square : a^2+b^2=53=25+10b+b^2+b^2 =>

=2b^2+10b-28=2(b^2+5b-14) |Viete's formula : sum=5, prod=-14 => -2*7)

=2(b-2)(b+7)

=> >The 2 numbers are (7,2), ( (-2,-7) is not solution since it should be positive)